OpenStudy (anonymous):

Solve: cos‏θ - (tan‏θcos‏θ)=0 for 0≤‏θ≤2pi

OpenStudy (anonymous):

\[\cos \Theta = \tan \Theta \cos \theta\]

OpenStudy (anonymous):

\[1 = \tan \theta\]

OpenStudy (anonymous):

\[\theta = \tan^{-1} (1)\]

OpenStudy (anonymous):

That's at pi/4, 5pi/4 between 0 and 2pi

OpenStudy (anonymous):

factorise cosΘ(tanΘ - 1) = 0 cos Θ =0, tanΘ = 1 Θ = pi/2 , pi/4 , 5pi/4